OpenStudy (anonymous):

what is t^2+12t-18 in vertex form? with steps please.

Nnesha (nnesha):

formula to find x-coordinate of the vertex \[\large\rm x=\frac{ -b }{ 2a}\] and then substitute x value into the equation to find y-coordinate Ax^2+Bx+C where a=leading coefficient and b=coefficient of x term c=constant

OpenStudy (anonymous):

i thought that equation gave you the slope?

Nnesha (nnesha):

which equation ??

OpenStudy (anonymous):

\[x=-b/2a\]

OpenStudy (yoloshroom):

(x2-x1)/(y2-y1 or smething gives you slope. this gives you the value of X, for vertex form you're looking for the formula of y=a(x-h)^2+k

Nnesha (nnesha):

no slope formula is different that's -b/2a to find x-coordinate of the vertex

Nnesha (nnesha):

or you can rewrite the equation in vertex form by completing the square

OpenStudy (anonymous):

how do you do that complete the square?

OpenStudy (yoloshroom):

explains it really well.

OpenStudy (anonymous):

i need to put t^2+12t-18 in vertex form by completing the square. i just reread the question.

Nnesha (nnesha):

oh so rewrite in vertex form i thought find the vertex

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

do you mean the x=-b/2a thing?

Nnesha (nnesha):

no well that's the formula to find x-coordinate of the vertex forget about it you need to rewrite the equation in vertex form by completing the square first step we need to get 1 for leading coefficient in order to complete the square

OpenStudy (anonymous):

what do you mean?

Nnesha (nnesha):

\[\large\rm y=t^2+12t-18\] in this equation a=1 so we can complete the square take half of x term and then add (b/2)^2 to both sides

Nnesha (nnesha):

what do you mean what ?

OpenStudy (anonymous):

so the new equation is .5t^2+6t-9?

Nnesha (nnesha):

no vertex form should be like this y=a(x-h)^2+k

OpenStudy (anonymous):

wait the x term is 6 so half that is 3... so the new equation is y=1(3-h)^2+k

OpenStudy (anonymous):

y=1(3(-6))^2+(-54)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

@Nnesha are you still there?

Nnesha (nnesha):

no sorry i was afk

Nnesha (nnesha):

how did you get -6 and we should keep the x variable

Nnesha (nnesha):

\[\large\rm y=t^2+12t-18\] in this equation a=1 so we can complete the square -18 is constant first we need to find perfect trinomail \[\rm y =(\color{Red}{t^2+12t}) -18\] now take half of x term and then square it (b/2)^2 add in the parentheses and subtract from constant term

Nnesha (nnesha):

let me know if you have a question so far